OpenStudy (anonymous):

(Calc Q > Derivatives > Chain Rule > Double chain rule scenarios?!) So I'm working on some problems for my homework and come across a problem that requires me to apply the chain rule twice. I have no clue where to begin, even though I do understand how to apply the chain rule. The problem in question is: g(x) = 2tan^3(4πx). The answer is 24π(tan^2(4πx))(sec^2(4πx). However having the answer helps me none since I do not know how to apply the chain rule twice... Help?! (MUCH thanks in advance!!!)

OpenStudy (kinggeorge):

Here's a step by step process.\[g(x) = 2\tan^3(4πx)\]\[{d \over dx}\quad 2\tan^3(4πx) \]\[=6\tan^2(4πx) \cdot {d \over dx}\;\;\tan(4\pi x)\]\[=6\tan^2(4πx) \cdot \sec^2(4 \pi x) \cdot {d \over dx}\;\;4\pi x\]\[=6\tan^2(4πx) \cdot \sec^2(4 \pi x) \cdot 4\pi\]\[=24 \pi \cdot \tan^2(4πx) \cdot \sec^2(4 \pi x)\]

OpenStudy (anonymous):

f(g(h(x)))' = f'(g(h(x))*g(h(x))' g(h(x))' = g'(h(x))*h'(x) So I get: f'(g(h(x))*g'(h(x))*h'(x)

OpenStudy (kinggeorge):

You basically just slowly move the \[d \over dx\]through the entire equation integrating with the chain rule one part at a time.

OpenStudy (anonymous):

Basically, just start on the outside function, take the derivative, and plug all of the inside functions into that derivative. Then multiply the result by the derivative of everything inside.

OpenStudy (anonymous):

erret, lets work on that... and while im trying this, could you come up with something else for me to try so i know that I get this?

OpenStudy (anonymous):

Yes. cos(sin(x^2)

OpenStudy (anonymous):

tan(4πx) comes from where? second part third step Would that be the inside of the original equasion?

OpenStudy (anonymous):

like what I don't understand is that usually, you would take d/dx(outside) x (inside) x d/dx(inside), as in d/dx(2(x^2-1)^2 ) = 4(x^2-1)(2x)

OpenStudy (anonymous):

Rewrite the function as: 2tan(4pix)^3 Chain rule says I can take the derivative as if it was just 2x^3, as long as I multiply by the derivative of the inside function after I do that. The inside in this case is tan(4pix)

OpenStudy (kinggeorge):

The first function you were integrating is \(u^3\) where \(u=\tan(4\pi x)\) So we take the derivative to get \[{d \over du} u^3=3u^2 \cdot \left({d \over du}\;\; u\right)\]By the definition of the chain rule. Now substitute \(u=\tan(4\pi x)\) back in, to get the expression that I wrote.

OpenStudy (anonymous):

This makes much more sense, now I shall attempt smoothmath's problem... btw thanks guys.

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (anonymous):

My pleasure =D

OpenStudy (anonymous):

alright so the problem, while somewhat similar to what ye gents described... well lol here is what I got? -sin(sinx^2)(cos(x^2)(2x)) I'm double-checking my work as I post, just wanted to get this out there

OpenStudy (kinggeorge):

That looks correct to me.

OpenStudy (anonymous):

spit out one more with slightly more complex numbers? :)

OpenStudy (kinggeorge):

How about this? \[-\cos(5\sin^2(x^4)+\tan(x^2))\]

OpenStudy (anonymous):

:D one moment please!

OpenStudy (anonymous):

I also agree with your answer for that first one.

OpenStudy (anonymous):

i'm going to try to apply the method used in http://www.youtube.com/watch?v=8qx5XsyXBQA , so i may take a little longer

OpenStudy (anonymous):

People who read cos as "cose" instead of "cosine" cringe...

OpenStudy (anonymous):

sin(5sin(x^8+tanx^2)(40sinx^7(cosx)+2(tanx)(sec^2x) is basically it, but i got careless with my parenthesis so i need to re-establish where the **** they need to be added to close the equasion -.-

OpenStudy (kinggeorge):

It looks close, but there's a few terms that are off. I certainly gave you a hard one.

OpenStudy (anonymous):

that would be agreeable. wolfram alpha also think that i am wrong, but I am a little lost on a few features on this answer.. >> http://www.wolframalpha.com/input/?i=d%2Fdx [%E2%88%92cos%285sin2%28x4%29%2Btan%28x2%29%29] << first things first, can 5sin^2(x^4) simplify to 5sin(x^4)^2)? and then must I double chain rule this or can I multiply in the ^2 to the ^4 to get 5sinx^8?

OpenStudy (kinggeorge):

You have to use the double chain rule on that.

OpenStudy (anonymous):

blehhhhhhhhhhhhhhhhhhhhhhhh lol okay. Well... thanks a ton for your help!! I think I am alright from here. :D

OpenStudy (kinggeorge):

You're very welcome. If you want a slightly easier problem to work, try \[-\cos(5\sin^2(x^4))\]Instead